OpenStudy (rizags):

Evaluate sin 18 degrees

OpenStudy (rizags):

using trig identities and formulas

OpenStudy (rizags):

yehp

OpenStudy (rizags):

i began with this: sin 18 = cos 72

OpenStudy (rizags):

not exact, i mean in terms of radicals

OpenStudy (rizags):

there is a way with manipulation of the trig identities

OpenStudy (rizags):

\[2\cos ^{2}36 - 1 = \cos 72\]

OpenStudy (rizags):

\[2(1-2\sin ^{2}18)^{2}-1 = \cos 72\]

OpenStudy (rizags):

then im lost

OpenStudy (rizags):

o ok

OpenStudy (mathmate):

@Rizags Are you familiar with the trigonometric products, example: \(\rm sin^2(x)=(1-cos(2x))/2\) If you apply it twice, you get the above identity, after which you can solve for sin(18)=cos(72)

OpenStudy (rizags):

so input x=36?